# Unified Multimodal Interleaved Document Representation for Retrieval

### v2

@@ 4 Experiments
Summary 13.4 41.3 66.5 21.6
Text-document 12.5 37.8 68.7 19.8
+ Single-image 16.4 45.4 77.1 25.3
+ Interleaved (Ours) 20.5 50.0 78.0 29.4
Table 2: Results with different section retrieval strategies. Document (Ours) performs document retrieval and section reranking, whereas Passage performs section retrieval and reranking. * denotes the model without reranking.
@@
Table 3: Performance on document retrievals. (a): Results of document retrieval for multimodal queries on InfoSeek and ViQuAE. (b): Results of document retrieval for textual queries on Encyclopedic-VQA (Enc-VQA) and ViQuAE.
(a) Document Retrieval with Multimodal Queries
Format Dataset R@1 R@10 R@100 MRR@10 Text-document InfoSeek 6.8 23.6 52.5 11.2 + Interleaved 10.2 30.4 57.3 15.7 Text-document ViQuAE 13.5 40.4 67.4 20.9 + Interleaved 17.5 46.0 69.4 26.3
(b) Document Retrieval with Textual Queries
Format Dataset R@1 R@10 R@100 MRR@10 Text-document Enc-VQA 62.7 76.3 87.4 67.0 + Interleaved 65.4 76.8 87.8 69.0 Text-document ViQuAE 55.8 71.5 83.0 60.9 + Interleaved 56.5 72.2 83.0 61.6
4.1 Experimental Setups
Datasets
We evaluate IDentIfy on four benchmark datasets designed for multimodal IR tasks that require understanding of both the textual and visual cues within queries and documents, as follows: Encyclopedic-VQA (Mensink et al., 2023) is a large-scale benchmark for multimodal Visual Question Answering (VQA) with queries linked to specific Wikipedia sections and includes both text-only and multimodal queries; InfoSeek (Chen et al., 2023) is a knowledge-intensive VQA dataset with multimodal questions generated from Wikidata triples that include diverse entities such as landmarks, animals, and food; ViQuAE (Lerner et al., 2022) involves both text-based and multimodal queries about human entities, linked to annotated Wikipedia sections, making it ideal for evaluating section retrieval; Open-WikiTable (Kweon et al., 2023) extends WikiSQL (Zhong et al., 2017) and WikiTableQuestions (Pasupat and Liang, 2015), targeting open-domain table QA by identifying documents or sections containing relevant tables. We provide more details on datasets in Appendix A.
Baselines
We compare our approach against diverse baselines that capture different document representations. First, the Entity and Summary baselines retrieve documents based on their titles and summary sections, respectively, leveraging high-level textual cues. Also, the Text-document retriever baseline utilizes the full textual content of documents for retrieval. We further include the Single-image baseline that additionally leverages the first image of each document. IDentIfy is our model that holistically represents multimodal content (text, images, and tables) within documents.
Evaluation Metrics
Implementation Details
We use LLaVA-NeXT-Interleave (Li et al., 2024b) of 0.5B parameters as the basis VLM for both the retriever and reranker. During training, documents are represented using randomly selected four sections, while in inference, we consider all sections within each document. For section-level retrieval, all sections within the top 25 retrieved documents are reranked. Experiments are conducted on a single H100 GPU.
4.2 Experimental Results and Analyses
@@
We report retrieval performance on the Encyclopedic-VQA dataset in Table 2, where queries include both text and images. We observe that IDentIfy achieves the best performance, improving R@1 scores by 53.0%, 64.0%, and 25.0% over Summary, Text-document, and Single-image retrieval baselines, respectively, with similar trends observed for other metrics. These results demonstrate the effectiveness of integrating multimodal content holistically into a unified representation. To further illustrate the advantages of our approach, we provide case studies in Appendix C.
We further examine the impact of our document retrieval and section reranking pipeline. In Table 2, the passage retriever represents individual sections as separate retrieval units, whereas the document retriever (ours) aggregates multiple section representations into a single representation. Then, we perform reranking over the retrieved sections or the sections from the retrieved documents, and then report the results in Table 2 (where * denotes the model without reranking). From this, we observe that the passage retriever without reranking (Passage*) achieves suboptimal retrieval performance, highlighting the challenge in pinpointing the most relevant section within a document using traditional retrieval methods. In contrast, when the reranker is used alongside the document retriever, the performance significantly surpasses that of the passage retrieval. These results confirm the importance of leveraging holistic context from multiple, interrelated sections within documents.
@@
Table 4: Performance on section reranking. (a): Results of section reranking for multimodal queries on Encyclopedic-VQA (Enc-VQA) and ViQuAE. (b): Results of section reranking for textual queries on Enc-VQA and ViQuAE. For fair comparison between different methods, the reranking target (i.e., candidate sections from retrieved documents) are the same across them.
(a) Section Reranking with Multimodal Queries
Format Dataset R@1 R@10 R@20 MRR@10 Text-document Enc-VQA 40.7 52.8 55.5 44.8 + Interleaved 42.4 53.6 55.7 46.3 Text-document ViQuAE 12.6 31.7 37.7 18.2 + Interleaved 11.4 32.1 39.2 17.5
(b) Section Reranking with Textual Queries
Format Dataset R@1 R@10 R@20 MRR@10 Text-document Enc-VQA 68.1 79.4 80.2 72.3 + Interleaved 69.7 80.1 80.6 73.6 Text-document ViQuAE 27.8 50.2 57.7 35.0 + Interleaved 29.9 50.9 59.8 36.7
Table 5: Retrieval results for tables, where Zero-shot denotes a model trained on Encyclopedic-VQA but not on the target dataset. Finetuned refers to additional training of the model on the target dataset. (a): Results for tabular document retrieval on Open-WikiTable (OWT). (b): Textual and tablular section retrieval results on ViQuAE and OWT datasets, respectively. (c): Reranker accuracy of a classification task that identifies the section containing the query-associated table given a gold document.
(a) Document Retrieval for Tables
Model R@1 R@10 R@100 MRR@10 Zero-shot 29.4 58.0 86.0 38.1 Finetuned 55.8 84.1 93.5 66.1
(c) Tabular Classification
Model Random Zero-shot Finetuned Acc@1 11.9 9.3 56.5
(b) Section Retrieval for Tables
Model Modality Dataset R@1 R@10 R@20 MRR@10 Zero-shot Text ViQuAE 20.3 49.0 57.7 28.9 Finetuned 29.9 50.9 59.8 36.7 Zero-shot Table OWT 5.9 20.5 29.4 9.1 Finetuned 8.4 36.7 52.8 15.2
We further expand our experiments to two additional datasets, InfoSeek and ViQuAE, and report document retrieval results. As shown in Table 3, our model consistently outperforms the Text-document baseline for both multimodal and text-only queries. We attribute these gains to the integration of multimodal content, allowing the VLM to capture richer alignments and leverage pre-existing knowledge for more effective document representation (Xu et al., 2024).
@@
Similarly, we evaluate section retrieval performance on Encyclopedic-VQA and ViQuAE datasets, for both multimodal and textual queries. As shown in Table 4, our model outperforms the Text-document baseline in most cases. However, the performance gains over the baseline are smaller compared to the document retrieval setup. This is likely because section reranking focuses on evaluating the relationship between a single section and a query (rather than leveraging the holistic context of the entire document), and individual sections may lack the diverse multimodal information necessary for fully capturing the intent of queries.
@@
We explore the retrieval task for tabular data, aiming to identify documents or sections containing query-relevant tables, and compare models trained on Encyclopedic-VQA (Zero-shot) with those additionally trained on Open-WikiTable (Finetuned). As shown in Table 5 (a), the Finetuned retriever outperforms the Zero-shot retriever on retrieving documents containing query-relevant tables. However, more fine-grained section reranking results (identifying sections containing query-relevant tables) in Table 5 (b) may reveal a notable modality-specific challenge: the performance of Zero-shot and Finetuned rerankers is considerably lower on table retrieval compared to their performance on text retrieval, despite both the text and tables being represented with word tokens. To better understand this, we design a classification task, where rerankers are tasked with identifying the correct section containing the target table within the golden document. Then, as shown in Table 5 (c), the Zero-shot reranker performs comparably to random selection, while the Finetuned reranker shows modest improvements. These findings highlight the intrinsic challenge of tabular retrieval, suggesting the need for table-specific modules to more holistically represent multimodal interleaved documents.
@@
Table 7: Comparison of different training objectives for the reranker: Contrastive considers sections as retrieval units and uses the one for document retriever training; Document + BCE concatenates the query with multiple sections from the same document and uses the BCL loss; Section + BCE trains the reranker by concatenating the query with each section individually.
(a) Section Retrieval for Multimodal Queries
Train Loss R@1 R@10 R@20 MRR@10 Contrastive 3.6 15.0 21.3 6.5 Document + BCE 13.6 29.6 32.9 24.1 Section + BCE (Ours) 42.4 53.6 55.7 46.3
(b) Section Retrieval for Textual Queries
Train Loss R@1 R@10 R@20 MRR@10 Contrastive 13.6 37.7 45.1 20.6 Document + BCE 23.8 43.4 47.2 39.1 Section + BCE (Ours) 69.7 80.1 80.6 73.6
To see how the number of sections used for representing each document impacts performance, we evaluate document retrieval on the InfoSeek dataset by varying the sections per document during training. As shown in Section 4.2, incorporating more sections improves MRR@10 from 7.5 to 15.7 due to leveraging richer multimodal and contextual information. However, this comes at the cost of increased computational requirements, as processing more sections raises GPU memory consumption.
In training the reranker, we investigate whether considering sections from the same document as negative examples (called In-document) is effective than other strategies, such as Top-K negatives (top-K retrieved sections based on their similarity with the input query) and In-batch negatives (positive sections from other samples in the same batch). As shown in Section 4.2, we observe that the In-document approach achieves superior performance especially on R@1, demonstrating its ability to effectively identify the most pertinent section among highly similar sections within the same document, i.e., its training objective can encourage the reranker to focus on fine-grained distinctions between closely related sections (within the same document).
BCE loss is the most effective to train the section reranker.
In our reranker design, we use a binary cross-entropy (BCE) loss by concatenating the query with each document section individually (Section + BCE), allowing the model to directly assess query-section relevance. As an alternative, we also explore a contrastive loss (Contrastive), which models section reranking similarly to document retrieval but uses sections as the retrieval units, and a variant of BCE loss (Document + BCE), where the query is concatenated with multiple sections (both positive and negative) from the same document. As shown in Table 7, the Section + BCE reranker outperforms both alternatives. Specifically, contrastive loss performs the worst, suggesting that direct concatenation of query and section provides clearer relevance signals, consistent with conventional reranking approaches. Moreover, while Document + BCE leverages inter-section context, its performance might be hindered by training constraints as the model processes fewer sections during training (Jiang et al., 2024a; Lee et al., 2024), and addressing it would be interesting future work.
Our Approach is Robust with Different Base Models.
Table 8: Results with another base model (LLaVA-OneVision) for document retrieval (with different document formats).
@@
Summary 7.6 24.7 55.7 12.0
Text-document 7.0 24.1 50.4 11.7
+ Single-image 9.3 31.4 61.9 15.4
+ Interleaved (Ours) 12.1 36.1 62.5 18.2
To ensure the effectiveness of our approach across VLMs, we evalulate its performance with another VLM, LLaVA-OneVision (Li et al., 2024a), with 0.5 billion paramters, in addition to LLaVA-NeXT-Interleave (Li et al., 2024b) used in our main experiments. Results in Table 8 show that ours continues to outperform baselines, achieving a notable 30.1% gain in R@1 over the best baseline.
@@ 5 Conclusion
In this paper, we introduced IDentIfy, a novel IR framework designed to address the limitations of conventional methods that rely on textual content of documents and their segmented passages. Specifically, our approach sits on top of recent VLMs, which enables integration and representation of diverse multimodal content (including text, images, and tables) into a unified document representation. Also, unlike previous strategies that segment documents at the passage level, our method merges these segments to maintain the document’s structural coherence, while further introducing a reranking strategy for precise identification of relevant sections. Extensive experiments across various IR datasets demonstrated that IDentIfy consistently outperforms existing baselines, confirming that the interleaved multimodal representation significantly enhances the quality of the document retrieval. We believe IDentIfy represents a crucial step toward more comprehensive and contextually aware IR systems, capable of handling the increasing multimodality of modern information sources.
@@
Due to the limitations of a single H100 GPU, we represent documents by selecting a limited number of sections and averaging their corresponding embeddings. While this reduces the computational demands, our findings suggest that capturing a broader document context leads to improved retrieval performance. Hence, leveraging the long context window of LVLMs could further enhance document retrieval by capturing more comprehensive information from the full document. Moreover, our reranker design follows the conventional approach of concatenating the input query with individual sections. However, we believe that providing the reranker with all the sections together would allow the model to better leverage the contextual information from the entire interleaved document, potentially resulting in improved performance. In order to fully leverage the interleaved format in the IR system, addressing the issues by reducing the GPU load when processing interleaved documents would greatly boost overall IR performance. We leave these explorations for future work.
Ethics Statement

### v3

@@ 4.1 Experimental Setups
Abstract 13.4 41.3 66.5 21.6
Text-Only 12.5 37.8 68.7 19.8
Text & Table 12.6 38.6 68.5 19.9
Text & Image 16.4 45.4 77.1 25.3
IDentIfy (Ours) 20.5 50.0 78.0 29.4
Table 1: Results with different document retrievers.
Evaluation Metrics
Implementation Details
We use LLaVA-NeXT-Interleave (Li et al., 2024b) as the basis VLM for both the retriever and reranker, and also use LLaVA-OneVision (Li et al., 2024a) as an additional basis VLM to show the robustness of IDentIfy. Following the convention of using the basis of retrieval with less than 1B parameters to balance computational efficiency and retrieval performance (Radford et al., 2021; Zhou et al., 2024; Wei et al., 2024), we choose 0.5B-parameter versions of the VLMs. During training, documents are represented using randomly selected four sections, while in inference, we consider all sections within each document. For section-level retrieval, all sections within the top 25 retrieved documents are reranked. Experiments are conducted on a single H100 GPU.
4.2 Experimental Results and Analyses
@@
Table 2: Comparison of different IR strategies for section retrieval. Document (Ours) performs document retrieval and section reranking, whereas Passage performs passage retrieval and reranking. * denotes the model without reranking.
@@
We report retrieval performance on the Encyclopedic-VQA dataset in Table˜1, where queries include both text and images. IDentIfy significantly outperforms all baselines built on VLMs such as CLIP and BLIP, which are limited to handling a single image alongside text and encoding image-text representations independently, making them suboptimal for understanding multimodal interactions within documents. We also observe that IDentIfy achieves the best performance, improving R@1 scores by 53.0%, 64.0%, 62.7%, and 25.0% over Abstract, Text-Only, Text & Table and Text & Image retrieval baselines, respectively, with similar trends observed for other metrics. These results demonstrate the effectiveness of integrating multimodal content holistically into a unified representation. To further illustrate the advantages of our approach, we provide case studies in Appendix˜E.
Document Retrieval Section Reranking
Dataset Query Type Method R@1 R@10 R@100 MRR@10 R@1 R@10 R@20 MRR@10
Enc-VQA Multimodal Text-Only 12.5 37.8 68.7 19.8 40.7 52.8 55.5 44.8
IDentIfy (Ours) 20.5 50.0 78.0 29.4 42.4 53.6 55.7 46.3
Textual Text-Only 62.7 76.3 87.4 67.0 68.1 79.4 80.2 72.3
IDentIfy (Ours) 65.4 76.8 87.8 69.0 69.7 80.1 80.6 73.6
ViQuAE Multimodal Text-Only 13.5 40.4 67.4 20.9 12.6 31.7 37.7 18.2
IDentIfy (Ours) 17.5 46.0 69.4 26.3 11.4 32.1 39.2 17.5
Textual Text-Only 55.8 71.5 83.0 60.9 27.8 50.2 57.7 35.0
IDentIfy (Ours) 56.5 72.2 83.0 61.6 29.9 50.9 59.8 36.7
InfoSeek Multimodal Text-Only 6.8 23.6 52.5 11.2 N/A N/A N/A N/A
IDentIfy (Ours) 10.2 30.4 57.3 15.7 N/A N/A N/A N/A
Table 3: Performance on document retrieval and section reranking for multimodal and textual queries on Encyclopedic-VQA (Enc-VQA), ViQuAE, and InfoSeek. We compare the approach that solely uses textual information from documents (Text-Only) and our approach of leveraging interleaved multimodal contents from the documents (IDentIfy) over various scenarios.
(a) Document Retrieval for Tables
Method R@1 R@10 R@100 MRR@10
Zero-shot 29.4 58.0 86.0 38.1
Finetuned 55.8 84.1 93.5 66.1
(c) Tabular Classification
Method Random Zero-shot Finetuned
Acc@1 11.9 9.3 56.5
(b) Section Reranking for Tables
Dataset Target Method R@1 R@10 R@20 MRR@10
ViQuAE Text Zero-shot 20.3 49.0 57.7 28.9
Finetuned 29.9 50.9 59.8 36.7
OWT Table Zero-shot 5.9 20.5 29.4 9.1
Finetuned 8.4 36.7 52.8 15.2
Table 4: Retrieval results for tables, where Zero-shot denotes a model trained on Encyclopedic-VQA but not on the target dataset. Finetuned refers to additional training of the model on the target dataset. (a): Results for tabular document retrieval on Open-WikiTable (OWT). (b): Textual and tablular section reranking results on ViQuAE and OWT datasets, respectively. (c): Reranker accuracy of a classification task that identifies the section containing the query-associated table given a gold document.
We further examine the impact of our pipeline of document retrieval and section reranking. In Table˜2, the passage retriever represents individual sections as separate retrieval units, whereas the document retriever (ours) aggregates multiple section representations into a single representation. Then, we perform reranking over the retrieved sections or the sections from the retrieved documents, and then report the results in Table˜2 (where * denotes the model without reranking). From this, we observe that the passage retriever without reranking (Passage*) achieves suboptimal retrieval performance, highlighting the challenge in pinpointing the most relevant section within a document using traditional retrieval methods. In contrast, when the reranker is used alongside the document retriever, the performance significantly surpasses the passage retrieval, demonstrating the effectiveness of our coarse-to-fine document-to-section retrieval strategy.
@@
We further expand our experiments to two additional datasets, InfoSeek and ViQuAE, and report document retrieval results. As shown in Table˜4 Left, our model consistently outperforms the Text-document baseline for both the multimodal and textual queries. We attribute these gains to the integration of multimodal content, allowing the VLM to capture richer alignments and leverage pre-existing knowledge for more effective document representation (Xu et al., 2024).
@@
Similarly, we evaluate section retrieval performance on Encyclopedic-VQA and ViQuAE datasets, for both multimodal and textual queries. As shown in Table˜4 Right, our model outperforms the Text-document baseline in most cases. However, the performance gains over the baseline are smaller compared to the document retrieval setup. This is likely because section reranking focuses on evaluating the relationship between a single section and a query (rather than leveraging the holistic context of the entire document), and individual sections may lack the diverse multimodal information necessary for fully capturing the intent of queries.
@@
Quesry Type Train Loss R@1 R@10 R@20 MRR@10
Multimodal Contrastive 3.6 15.0 21.3 6.5
Doc + BCE 13.6 29.6 32.9 24.1
Sec + BCE (Ours) 42.4 53.6 55.7 46.3
Textual Contrastive 13.6 37.7 45.1 20.6
Doc + BCE 23.8 43.4 47.2 39.1
Sec + BCE (Ours) 69.7 80.1 80.6 73.6
Table 5: Comparison of training objectives for the reranker: Contrastive uses contrastive loss similar to the document retriever training; Doc + BCE concatenates the query with multiple sections from the same document and uses the BCE loss; Sec + BCE trains the reranker by concatenating the query with each section individually.
@@
Abstract 7.6 24.7 55.7 12.0
Text-Only 7.0 24.1 50.4 11.7
Text & Table 6.9 26.3 54.9 12.1
Text & Image 9.3 31.4 61.9 15.4
IDentIfy (Ours) 12.1 36.1 62.5 18.2
Table 7: Results with another base model (LLaVA-OneVision-0.5B) for document retrieval (with different document formats).
We explore the retrieval task for tabular data, aiming to identify documents or sections containing query-relevant tables, and compare models trained on Encyclopedic-VQA (Zero-shot) with those additionally trained on Open-WikiTable (Finetuned). As shown in Table˜4 (a), the Finetuned retriever outperforms the Zero-shot retriever on retrieving documents containing query-relevant tables. However, more fine-grained section reranking results (identifying sections containing query-relevant tables) in Table˜4 (b) may reveal a notable modality-specific challenge: the performance of Zero-shot and Finetuned rerankers is considerably lower on table retrieval compared to their performance on text retrieval, despite both the text and tables being represented with word tokens. To better understand this, we design a classification task, where rerankers are tasked with identifying the correct section containing the target table within the golden document. Then, as shown in Table˜4 (c), the Zero-shot reranker performs comparably to random selection, while the Finetuned reranker shows modest improvements. These findings highlight the intrinsic challenge of tabular retrieval, suggesting the need for table-specific modules to more holistically represent multimodal interleaved documents.
To see how the number of sections used for representing each document impacts performance, we evaluate document retrieval on the InfoSeek dataset by varying the sections per document during training. As shown in Figure˜3, incorporating more sections improves MRR@10 from 7.5 to 15.7 due to leveraging richer multimodal and contextual information. However, this comes at the cost of increased computational requirements, as processing more sections raises GPU memory consumption.
BCE loss is the most effective to train the section reranker.
In our reranker design, we use a binary cross-entropy (BCE) loss by concatenating the query with each document section individually (Section + BCE), allowing the model to directly assess query-section relevance. As an alternative, we also explore a contrastive loss (Contrastive), which models section reranking similarly to document retrieval but uses sections as the retrieval units, and a variant of BCE loss (Document + BCE), where the query is concatenated with multiple sections (both positive and negative) from the same document. As shown in Table˜5, the Section + BCE reranker outperforms both alternatives. Specifically, contrastive loss performs the worst, suggesting that direct concatenation of query and section provides clearer relevance signals, consistent with conventional reranking approaches. Moreover, while Document + BCE leverages inter-section context, its performance might be hindered by training constraints as the model processes fewer sections during training (Jiang et al., 2024b; Lee et al., 2024), and addressing it would be interesting future work.
In training the reranker, we investigate whether considering sections from the same document as negative examples (called In-document) is effective than other strategies, such as Top-K negatives (top-K retrieved sections based on their similarity with the input query) and In-batch negatives (positive sections from other samples in the same batch). As shown in Table˜7, we observe that the In-document approach achieves superior performance especially on R@1, demonstrating its ability to effectively identify the most pertinent section among highly similar sections within the same document, i.e., its training objective can encourage the reranker to focus on fine-grained distinctions between closely related sections (within the same document).
Our IDentIfy is Versatile with Different VLMs.
To ensure the effectiveness and robustness of IDentIfy across VLMs, we evalulate its performance with another VLM, LLaVA-OneVision (Li et al., 2024a), with 0.5 billion parameters, in addition to LLaVA-NeXT-Interleave (Li et al., 2024b) used in our main experiments. Results in Table˜7 show that ours continues to outperform baselines, achieving a notable 30.1% gain in R@1 over the best baseline.
@@ 5 Conclusion
In this paper, we introduced IDentIfy, a novel IR framework designed to address the limitations of conventional methods that rely on textual content of documents and their segmented passages. Our approach sits on top of recent VLMs, which enables integration and representation of diverse multimodal content (including text, images, and tables) into a unified document representation. Also, unlike prior strategies that segment documents at the passage level, our method merges these segments to maintain the document’s structural coherence, while further introducing a reranking strategy for precise identification of relevant sections. Extensive experiments across various IR datasets show that IDentIfy consistently outperforms baselines, confirming the value of interleaved multimodal representation for both document and section retrieval. We believe IDentIfy represents a crucial step toward more comprehensive and contextually aware IR systems, capable of handling the increasing multimodality of modern information sources.
@@
Due to the constraints of a single H100 GPU that we have, we represent documents by sampling a limited number of sections and averaging their corresponding embeddings (See Figure˜3). While this reduces the computational demands, our findings suggest that capturing a broader document context leads to improved retrieval performance. Hence, leveraging the long context window of LVLMs with a greater number of sections could further enhance document retrieval by capturing more comprehensive information within the full document. Additionally, while using the basis model size of 0.5B (or less than 1B) parameters is a standard practice in IR literature, scaling up the basis VLMs remains an avenue for future work; however, although larger models can yield performance gains, they come at the cost of increased computational requirements. Moreover, our reranker design follows the conventional approach of concatenating the input query with individual sections. However, we believe providing the reranker with all sections together would allow the model to better leverage the contextual information from the entire document, potentially resulting in improved performance, and we leave explorations on this for future work.
Ethics Considerations
@@
Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00256259), the grant of the Korea Machine Learning Ledger Orchestration for Drug Discovery Project (K-MELLODDY) funded by the Ministry of Health & Welfare and Ministry of Science and ICT, Republic of Korea (grant number: RS-2024-12345678), the InnoCORE program of the Ministry of Science and ICT (N10250156), and the Institute for Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program (KAIST)).